# User Simulation for Evaluating Information Access Systems

### v1

@@ 7.1.1 Dialogue Acts
A taxonomy of actions designed specifically for conversational search is presented in [Azzopardi:2018:CAIR], with reference to the high-level desiderata in [Radlinski:2017:CHIIR] as well as the different phases of the search process in [Trippas:2018:CHIIR]; see Table 7.1. Overall, it is assumed that the conversational process revolves around the system trying to help the user resolve their information need. There are two main types of user actions, which either (1) change the state of the information need or (2) are related to the space of information objects. The first type of actions (Reveal) corresponds to the user disclosing details regarding their information need, which may also evolve during the search. The second type of actions involves inquiring about the available options (Inquire), navigating them (Navigate), or asking the system how it understood the information need or the why it presented particular suggestions (Interrogate). Additionally, there may be points in the conversation when the user wants to interrupt or stop the system (Interrupt). Azzopardi:2018:CAIR note that their list attempts to represent the main actions previously observed and discussed in the literature, but it is non-exhaustive and is meant to be taken as a starting point. For example, Bernard:2023:SIGIR expand a selected set of communicative functions from ISO 24617-2 with intents from [Azzopardi:2018:CAIR] in order to characterize multi-goal conversations in an e-commerce setting.
7.1.2 Dialogue Structure
@@
Dialogue structure can be designed by envisioning the optimal sequencing of turns, identifying key dialogue acts and their dependencies, anticipating potential user intents and system responses, and incorporating strategies for effective information flow. For example, Zhang:2018:CIKM suggest to classify interactions in an e-commerce conversational search setting into three stages: initiation, conversation, and display. Aliannejadi:2021:CIKM propose a user model for mixed-initiative conversational search that consists of three major phases: querying, feedback, and browsing. Zhang:2020:KDD propose a dialogue workflow model (CIR6) specific to the task of conversational recommendation by defining six key user intents and their dependencies; see Fig. 7.1.
@@
Another way of identifying structural patterns in dialogues is through the annotation and analysis of transcripts of human-human conversations, which represent an ideal setting in terms of how people naturally interact. For example, Qu:2018:SIGIR find that a frequent user intent transition pattern in information-seeking conversations is: START $\Rightarrow$ original question ($\Rightarrow$ potential answer $\Rightarrow$ further details)x3 $\Rightarrow$ potential answer $\Rightarrow$ positive feedback $\Rightarrow$ END. Lyu:2021:WWW analyze context-driven recommendation in the restaurant domain and break conversations down into three distinct stages, based on the recommendations made (0, 1, or 2+). They observe that conversations shift from preference elicitation and refinement in the first stage to inquiry and critiquing in subsequent stages, with additional comparisons in the last stage. A similar pattern is identified in [Bernard:2023:SIGIR], analyzing multi-goal dialogues in an e-commerce setting: the first part of the conversations tends to focus on uncovering the user’s needs and preferences to recommend some products (i.e., recommendation), while the second part is mostly concerned with product-related information seeking (i.e., search and QA), with users sometimes asking for other options (i.e., a second peak in recommendation). Zhang:2022b:SIGIR study how users reformulate their requests in case a conversational recommender system fails to understand them. They observe a typical pattern where a persistent user would first try to rephrase, then simplify before giving up. They also observe that rephrasing tends to happen earlier in dialogues than simplification.
@@
Similar to the workflows of information seeking that we discussed earlier (see Section 6.1), the dialogue structures provide a basis for defining the states in the MDP framework for user simulation. Modeling the transitions between different states would be necessary for simulating a user’s actions in the right context of the conversation.
7.2 Simulator Architectures
The main challenge in simulating users in a conversational setting is that the simulator needs to have advanced natural language understanding and generation capabilities, much like that of a conversational system. Hence, adopting the architecture of task-oriented dialogue systems is a natural approach in the construction of user simulators. Accordingly, two broad categories of simulator architectures can be distinguished: modular (or pipelined) and end-to-end systems. The difference between these two is more fundamental than just an architectural choice, it is also a question of granularity: modular simulators model user responses semantically on the level of dialogue acts, and then generate the corresponding natural language utterances. End-to-end systems, on the other hand, operate on the utterance level, i.e., generate textual responses directly. An analogy may be drawn between these two architectures and the classes of model-based vs. data-driven simulation approaches (cf. Section 1.4.1). While end-to-end systems might yield more fluent dialogues, they do not allow us to interpret user behaviour, and therefore are not that well suited for our purposes. In particular, with an end-to-end system, it is difficult or impossible to vary any trained simulator to simulate variable user behaviour. Parameterization of such systems with meaningful user variables would be desirable in general, but this requires modeling a user beyond their surface-level utterances in order to capture latent user states using more interpretable models.
7.2.1 Modular Systems
Dialogue management is responsible for maintaining the dialogue state and determining the next user action based on the system’s utterance and the goals, knowledge, and preferences of the individual. The dialogue state in traditional task-oriented dialogue systems is based around the notion of a semantic frame, which is a knowledge structure consisting of a collection of slots that together specify what the system needs to know to complete a given task [Jurafsky:2023:book]. [^2] In spoken dialogue systems, a considerable research effort is centered on dealing with the uncertainty resulting from noise in automatic speech recognition. There, it is common for the dialogue system to maintain multiple hypotheses of the true dialog state [Williams:2016:DD]. However, for text-based interactions, this is typically not a concern. For user simulation, the dialogue state needs to capture dialogue context from the user’s perspective. This includes representing the current information need, the dialogue history, and the system’s last action. The dialogue state may be represented explicitly, in a directly observable and interpretable way, such as previous dialogue acts and the status of various slots, or as a latent representation in an embedding space. The dialogue policy determines how the user should respond, i.e., the user intent and associated parameters; this might be influenced by the individual characteristics of the user. To ensure that the simulator maintains a dialogue strategy, a commonly used approach is an agenda-based simulator; it maintains an agenda or a stack of actions and determines which action to execute at each turn. Alternatively, the policy may be learned directly from data. We discuss these alternatives in more detail in Section 7.3.
Natural language generation (NLG) is concerned with turning the generated response from a structured representation (dialogue act) into natural language. At its simplest form, this is based on templates [Gur:2018:SLT, Shi:2019:EMNLP]. Other alternatives include retrieval-based methods [Wu:2017:ACL, Hu:2014:NIPS], neural text generation models [Wen:2015:EMNLP, Tseng:2021:ACL], and hybrid solutions [Li:2017:arXiv]. Natural language generation can also consult the user modeling component to produce outputs that are customized based on individual characteristics.
User modeling focuses on capturing the characteristics of individuals that would influence how they interact with the system. These could include, among others, information about the user’s goal, knowledge, preferences, personal characteristics (e.g., patience), and beliefs about the system (e.g., based on past interactions). In case of information access tasks, one can often identify commonalities (similar user state variables) between user modeling in a conversational system and traditional search/recommender systems. However, both the information available for inferring a user’s state and the observable actions are different. In particular, for conversational systems, we would need to model how a user’s latent state influences the natural language utterance generated by the user as well as infer a user’s latent state based on the observed natural language conversations (when training a simulator).
7.2.2 End-to-end Systems
Figure 7.4: End-to-end user simulator.
More recently, there have been numerous attempts to build end-to-end trainable task-oriented dialogue systems, leveraging advances in end-to-end neural generative models [Gao:2019:FnTIR, Ni:2022:AIR]. The main benefits over modular systems include the possibility of global optimization and easier adaptation to new domains [Chen:2017:SIGKDD]. End-to-end architectures may also be used for constructing user simulators, by casting the generation of user responses as a sequence-to-sequence (or source-to-target transduction) problem: given a dialogue history up to the given turn as input, generate the user’s natural language response as the output sequence. [^3] As noted above for modular architectures, in end-to-end simulation the NLU component may also be omitted [Lin:2022:SIGDIAL]. Instead, the dialogue act corresponding to the system utterance as taken as input, effectively equipping the simulated user with perfect language understanding capabilities. This problem can be effectively addressed by transformer-based natural language generation models [Tseng:2021:ACL, Lin:2022:SIGDIAL]. However, as indicated earlier, such model-based end-to-end approaches are more suitable for other uses of simulation, such as RL-based learning [Lin:2022:SIGDIAL, Tseng:2021:ACL] or data augmentation [Wan:2022:EMNLP, Mohapatra:2021:EMNLP, Kim:2021:ACL]. For simulation-based evaluation, a more controllable generation of user responses is required, which is (currently) cannot be guaranteed with end-to-end systems.
7.3 User Dialogue Policy
Assuming a modular simulator architecture, the user’s response is governed by a dialogue policy (a.k.a. dialogue strategy). In this section, we discuss work on modeling user policy in the context of task-oriented dialogue systems. It is worth emphasizing that all of the studies covered here treat task-oriented dialogue in a restricted “slot-filling” sense. Specifically, systems are built for specific domains, each characterized by a domain ontology that describes the specific intents, slots, and entities that can be talked about. The supported conversational functionality is limited to the user specifying their constraints in terms of informable slots and requesting information on requestable slots to find matching entities in a knowledge base. This matches some of the scenarios that are being considered within conversational information access, such as finding items where the user already has a reasonably concrete idea of attributes and the constraints on them, or looking up known items or their properties. However, CIA systems ultimately support a much broader array of information needs. In the case of conversational item recommendation, for example, the goal can naturally be characterized in terms of slot-value pairs [Zhang:2020:KDD]. However, for information-seeking or more exploratory scenarios the representation of user goal is an open research problem.
Formally, human-machine dialogue can represented semantically as a sequence dialogue acts by the system ($a_{i}^{s}$) and the user ($a_{i}^{u}$) as they take turns: $a_{0}^{s}\rightarrow a_{0}^{u}\rightarrow a_{1}^{s}\rightarrow a_{1}^{u}\rightarrow\dots\rightarrow a_{t-1}^{s}\rightarrow a_{t}^{s}$. A dialogue act is a tuple $a=(I,V)$, where $I\in\mathcal{I}$ is an intent from the set of possible dialogue intents (defined separately for user $\mathcal{I}^{u}$ and system $\mathcal{I}^{s}$) and $V$ is a list of slot-value pairs $V=\langle(slot_{1},value_{1}),\dots,(slot_{n},value_{n})\rangle$. The policy $\pi$ determines what action $a^{u}_{t+1}$ the user should take next, given the dialogue history. A main challenge here is that the simulator should give responses that are coherent across multiple conversational turns and are driven by some user goal. This formal framework can be connected with the general MDP framework that we discussed in the previous chapter by viewing the dialogue history as being represented as part of the state in the MDP framework.
Broadly speaking, approaches can be categorized into three levels: handcrafted methods that rely on a limited number of parameters and basic rules, models that are learned entirely from data, and hybrid solutions that fall somewhere in between. It is also interesting to observe that while early work was done exclusively in the field of dialogue systems [Schatzmann:2006:KER], more recent work is being increasingly carried out within the field of natural language processing. Below, we present some of the most influential approaches, categorized into statistical user models and sequence-to-sequence models. Within each category, approaches are discussed in a roughly chronological order, highlighting the progression of methods and how subsequent methods build upon and are influenced by their predecessors.
7.3.1 Statistical User Models
Statistical user modeling techniques “exploit the Markov assumption that the next user action can be predicted based on some representation of the current state of the dialogue” [Schatzmann:2006:KER]. As the user and system take turns, the simulator receives the corresponding dialogue acts ($a^{u}_{t}$ for user and $a^{s}_{t}$ for system), and updates the state:
N-grams Models
$\pi(s_{t})=P(a^{u}_{t+1}|a^{s}_{t})~{},$ (7.3)
where the conditional probabilities are estimated from an annotated corpus of human-machine dialogues. This model does not have any information about the user’s goal and does not place any constraints on the simulated user behaviour, therefore fails to produce realistic dialogues. Subsequent work tried to correct for it by estimating a selected set of simpler probabilities corresponding to “sensible” pairs of system and user actions (specifically, responding to greetings, constraining questions, and relaxation prompts), while other are assumed to be zero, thereby setting some constraints on the flow of the dialogue [Levin:2000:TSAP]. [^4] In principle, all the selected probabilities could be estimated from a dialogue corpus, but in practice, the probabilities corresponding to constraining or relaxing questions are set manually in [Levin:2000:TSAP] as they were not present in the corpus used. While this yields somewhat more realistic dialogues, the consistency between user responses across the dialogue is still not guaranteed. Also, simulated dialogues might run infinitely long if the user keeps changing their answer to the same question.
Goal-directed User Model with Memory
Agenda-based Simulator
In [Schatzmann:2007:NAACL] the agenda is initialized with all goal constraints set to INFORM acts and all goal requests set to REQUEST acts; the BYE act is added at the bottom of the agenda to close the dialogue. For the above example, this corresponds to:
$P(s_{t+1}|s_{t},a_{t}^{u})$ $=P(A_{t+1},G_{t+1}|A_{t},G_{t},a_{t}^{u})$ (7.5)
$=\delta(A_{t+1},A_{t}[1..N_{t}])\>\delta(G_{t+1},G_{t})~{}.$
In Schatzmann:2007:NAACL, the agenda update ($A_{t}\rightarrow A_{t+1}$) is viewed as a sequence of push operations, where dialogue acts get added on top of the agenda. This is typically performed using handcrafted heuristics. For example, if a $slot_{i}=value_{i}$ provided in $a^{s}_{t}$ violates one of the goal constraints in $G_{t}$, then one of the following may be pushed onto $A_{t+1}$: NEGATE, INFORM($slot_{i}=value_{j}$), or DENY($slot_{i}=value_{i},slot_{i}=value_{j}$). The push operations are also followed by a deterministic “clean up” step, where duplicated or unnecessary dialogue acts (e.g., requests for already filled goals) are removed.
Zhang:2020:KDD present an agenda-based user simulator for evaluating conversational recommender systems. They perform agenda initialization and update operations based on probabilities learned from a sample of annotated dialogues, with the help of a task-specific interaction model (cf. Fig. 7.1). There, state transitions are restricted to a prescribed number of transitions to other states based on user intents. Additionally, each user intent is mapped to a set of system intents that are considered appropriate responses (e.g., an INQUIRE user action can accept LIST or ELICIT as a system response). This allows agenda updates to conditioned on whether the system responds “as expected,” and the component probabilities are simple enough to be estimated directly from training data.
7.3.2 Sequence-to-sequence Models
In a more recent line of research, sequence-to-sequence models have been employed to create user simulators that can be learned in a fully data-driven manner from dialogue corpora; Table 7.2 presents a high-level overview.
Table 7.2: Overview of user simulators using a sequence-to-sequence model structure.
 Reference Architecture Input Output Modeling Multi-
goal? domain?
 [ElAsri:2016:Interspeech] RNN-LSTM feature vect. dialogue act Y N
 [Gur:2018:SLT] RNN-GRU dialogue act dialogue act Y N
 [Lin:2021:SIGDIAL] Transformer feature vect. dialogue act Y Y
 [Crook:2017:Interspeech] RNN-GRU/LSTM utterance utterance N N
 [Kreyssig:2018:SIGDIAL] RNN-LSTM feature vect. utterance Y N
 [Lin:2022:SIGDIAL] Transformer ${}^{\dagger}$ dial. act + utt. Y Y
${}^{\dagger}$ Last system dialogue act, previous three user dialogue acts, user goal, and turn number as JSON-formatted string
ElAsri:2016:Interspeech propose an LSTM-based approach that takes as input a sequence of dialogue contexts and outputs a sequence of user actions. At the beginning of each dialogue, a user goal with constraints and requests, $G=(C,R)$, is randomly generated, similar to the agenda-based approach, which remain fixed during the dialogue. In each turn $t$, the simulated user takes as input the conversation contexts observed so far, $\langle c_{1},\dots,c_{t}\rangle$. Specifically, $c_{t}$ consists of four components (all represented as binary vectors): the most recent machine action, the inconsistency between machine information and user goal (i.e., slots that have been misunderstood by the system so that these may be corrected), the constraint status (to inform the system about preferences), and the request status (to keep track of requests that have not yet been fulfilled). This is passed to an RNN, which encodes this dialogue history into a single vector $v_{t}$. Then, $v_{t}$ is used to initialize the decoder RNN at each time step, which outputs a sequence of $l$ user intents $\langle I_{1},\dots,I_{l}\rangle$. These intents are turned to dialogue acts $a_{t+1}^{u}$ by looking at the current user goal and uniformly drawing among the constraints/requests left. Gur:2018:SLT propose an end-to-end hierarchical user simulator (HUS), which encodes the entire dialogue history based on the user goal and system dialogue act, without requiring feature extraction or a supervised signal for state tracking. A pivotal aspect of their approach is to represent system dialogue acts on a more coarse level, by replacing specific slot values with one of the following: Requested, if the value is requested by the system, ValueInGoal, if the value appears in the user goal, ValueContradictsGoal, if the value contradicts the user goal, DontCare, if the value in the user goal is flexible, or Other otherwise. The actual values can then be substituted either from the user goal, system dialogue act, or from the knowledge base. A modified version of their model, the variational hierarchical user simulator (VHUS), introduces a variational sampling step before decoding the user turn to generate more diverse responses. Gur:2018:SLT also develop a goal regularization approach to enforce a more direct correspondence and avoid generating too long dialogues when the user diverges from the initial goal. Lin:2021:SIGDIAL present a transformer-based user simulator (TUS), which employs a similar feature-based input representation to [ElAsri:2016:Interspeech], but design it to be domain-independent. They use variable input lengths based on the slots mentioned by the system. Using a transformer model as the user policy network, input sequences of arbitrary lengths are handled and the relationships between the different slots can be captured by self-attention. The output includes the domain that should be selected for the given turn and a list of one-hot vectors to determine the values of the slots, using a coarse-grained representation similar to [Gur:2018:SLT]. (Note that the user intent is not explicitly predicted, but since only two possible user intents are considered, INFORM and REQUEST, those can be inferred based on how slot-values are represented.)
Unlike the above model that works on the semantic level of dialogue acts, sequence-to-sequence models have also been used to generate the text of user utterances directly. Crook:2017:Interspeech extend a basic encoder-decoder neural network with an additional dense layer between the encoder and decoder for better representation of context, following a similar architecture that has been proposed for non-goal-driven dialogue systems in [Serban:2016:AAAI]. Kreyssig:2018:SIGDIAL present the Neural User Simulator (NUS), which follows a very similar architecture and context representation as [ElAsri:2016:Interspeech], but the output is a sequence of words. NUS allows for goal changes (when the system indicates that the goal is unachievable) and is able to generate a diverse set of utterances in the same dialogue context.
Most recently, Lin:2022:SIGDIAL present a generative transformer-based simulator (GenTUS) that generates semantic as well as natural language representations of the user action, that is, both a dialogue act and an utterance. This approach harnesses the capabilities of pre-trained large language models, which helps with generating fluent language and with generalization to new domains, but may suffer from hallucinations. Specifically, GenTUS takes the last system dialogue act, the previous three user dialogue acts, the user goal, and the turn number $t$ as input. The decoder generates the next user dialogue act and the corresponding language response as output, leveraging the ontology to constrain the generation process to prevent generating illegal semantic actions.
In conclusion, a noticeable trend can be observed in model development (cf. Table 7.2), transitioning from manual feature engineering to progressively adopting end-to-end approaches. As these models become more end-to-end, their interpretability diminishes, while offering limited control over the behaviour of the simulated user. Effectively, dialog planning can only be controlled indirectly, through the input data that is supplied for training, which is problematic for several reasons. First, there is sufficient amount of training data required in the target domains, which makes these approaches less effective in low-resource scenarios. Second, if some actions or conversation paths are not present in the training corpus then those will not be captured in the simulator. Third, data-driven simulators may suffer from bias present in the corpus. Fourth, if users with various personalities are to be modeled (e.g., based on the level of cooperativeness), these characteristics need to be provided as supervised training signals along with each training example [Gur:2018:SLT]. Different experiments might require additional data annotation/collection effort, which would contradict the primary goal of using simulation as a means of automatic evaluation without human involvement. Generally, basing user dialogue policy on sequence-to-sequence models suffers from the same fundamental issues as end-to-end simulators; cf. our earlier discussion in Section 7.2.2.
7.4 User Simulation for Conversational Information Access
In this section, we present approaches for simulating user utterances that have been developed specifically for conversational information access. Conversational information access is a broad task that encompasses the goals of conversational search, recommendation, and question answering. However, approaches that support this holistic view are yet to be developed. It is only very recently that datasets that support multiple conversational goals have started to become available [Bernard:2023:SIGIR]. While it can be expected that multi-goal simulators will follow, there is no such published work at the time of writing. In existing work, simulation approaches are developed in a goal-specific manner, therefore we discuss them along this distinction: conversational search (Section 7.4.1) and conversational recommendation (Section 7.4.2).
7.4.1 Conversational Search
Conversational search corresponds to a special type of task-oriented dialogue, where the goal is to satisfy the user’s information need. Broadly speaking, there are two main types of user utterances considered: user-initiated questions and responses to system-initiated questions.
Simulating User Questions

### v2

@@ 7.4.1 Action Space
A taxonomy of actions designed specifically for conversational search is presented in [Azzopardi:2018:CAIR], with reference to the high-level desiderata in [Radlinski:2017:CHIIR] as well as the different phases of the search process in [Trippas:2018:CHIIR]; see Table 7.2. Overall, it is assumed that the conversational process revolves around the system trying to help the user resolve their information need. There are two main types of user actions, which either (1) change the state of the information need or (2) are related to the space of information objects. The first type of actions (Reveal) corresponds to the user disclosing details regarding their information need, which may also evolve during the search. The second type of actions involves inquiring about the available options (Inquire), navigating them (Navigate), or asking the system how it understood the information need or why it presented particular suggestions (Interrogate). Additionally, there may be points in the conversation when the user wants to interrupt or stop the system (Interrupt). Azzopardi:2018:CAIR note that their list attempts to represent the main actions previously observed and discussed in the literature, but it is non-exhaustive and is meant to be taken as a starting point. For example, Bernard:2023:SIGIR expand a selected set of communicative functions from ISO 24617-2 with intents from [Azzopardi:2018:CAIR] in order to characterize multi-goal conversations in an e-commerce setting.
Dialogue Structure
@@
Dialogue structure can be designed by envisioning the optimal sequencing of turns, identifying key dialogue acts and their dependencies, anticipating potential user intents and system responses, and incorporating strategies for effective information flow. Trippas:2018:CHIIR provide a different high-level formalization, distinguishing three stages of the conversational search process: (1) query formulation, (2) search results exploration, and (3) query re-formulation. Zhang:2018:CIKM suggest to classify interactions in an e-commerce conversational search setting into three stages: initiation, conversation, and display. Aliannejadi:2021:CIKM propose a user model for mixed-initiative conversational search that consists of three major phases: querying, feedback, and browsing. Zhang:2020:KDD propose a dialogue workflow model (CIR6) specific to the task of conversational recommendation by defining six key user intents and their dependencies; see Fig. 7.3.
@@
Another way of identifying structural patterns in dialogues is through the annotation and analysis of transcripts of human-human conversations, which represent an ideal setting in terms of how people naturally interact. Qu:2018:SIGIR find that a frequent user intent transition pattern in information-seeking conversations is: START $\Rightarrow$ original question ($\Rightarrow$ potential answer $\Rightarrow$ further details)x3 $\Rightarrow$ potential answer $\Rightarrow$ positive feedback $\Rightarrow$ END. Lyu:2021:WWW analyze context-driven recommendation in the restaurant domain and break conversations down into three distinct stages, based on the recommendations made (0, 1, or 2+). They observe that conversations shift from preference elicitation and refinement in the first stage to inquiry and critiquing in subsequent stages, with additional comparisons in the last stage. A similar pattern is identified in [Bernard:2023:SIGIR], analyzing multi-goal dialogues in an e-commerce setting: the first part of the conversations tends to focus on uncovering the user’s needs and preferences to recommend some products (i.e., recommendation), while the second part is mostly concerned with product-related information seeking (i.e., search and QA), with users sometimes asking for other options (i.e., a second peak in recommendation). Zhang:2022b:SIGIR study how users reformulate their requests in case a conversational recommender system fails to understand them. They observe a typical pattern where a persistent user would first try to rephrase, then simplify before giving up. They also observe that rephrasing tends to happen earlier in dialogues than simplification.
@@
Similar to the information-seeking workflows discussed earlier (see Section 6.1), dialogue structures can provide a basis for defining the states within the MDP framework for user simulation. Modeling the transitions between states would be necessary for simulating a user’s actions in the right context of the conversation To enhance the realism of simulations, we can draw upon extensive past work from linguistics and psychology on how conversations are structured and their common patterns, ranging from immediate “adjacency pairs” [Sacks:1974:L] to higher-level discourse patterns [Thomas:2021:TOIS].
7.5 User Simulation for Conversational Search
Conversational search is a type of task-oriented dialogue, where the goal is to satisfy the user’s information need. We have discussed in Section 7.4.1 the space of user actions for information-seeking conversations. This, combined with an appropriate state representation, could allow for modeling the user’s decision-making process as an MDP. However, a key challenge lies in creating a computational representation of the user’s goal and modeling progress towards the completion of that goal as part of the user’s state. When system responses are not limited to predefined information objects (e.g., passages or documents) but can be arbitrary generated text, relying on preexisting object-level relevance judgments is no longer feasible. Consequently, current user simulation techniques often make strong assumptions about the dialogue structure and require responses to be documents/paragraphs with existing relevance judgments. Given that mixed-initiative interaction is a defining characteristic of conversational information access, much of the research has focused on simulating responses to clarification questions. Notably, current research is limited to simulating specific user actions within a conversational search dialogue, with no existing methods capable of fully automatic, end-to-end dialogue simulation.
In the following, we discuss existing approaches categorized according to being model-based or data-driven.
7.5.1 Model-Based Approaches